# The James Webb Interferometer: Space-based interferometric detections of PDS 70 b and c at 4.8 m

### v1

@@ 5.2 Contrast Limits and the Nature of the Signal Seen in the Residuals
The residual signal shown in Figure 5, is in a direction significantly offset from the expected forward scattering direction of the inner disk. This indicates that what we observe is not due to a simple inner disk structure, and may hint at a complex inner disk morphology such as a spiral or clumpy features, as has been suggested by Casassus & Cárcamo (2022), from ALMA observations in the mm, and by Mesa et al. (2019) from near-infrared observations with VLT/SPHERE. It is also possible that the asymmetry we observe is related to the larger separation, 180 mas, ALMA compact emission seen by Balsalobre-Ruza et al. (2023), as it is at very similar position angle. It might even highlight extended emission from an accretion stream between PDS 70 b and c, tentatively detected by Christiaens et al. (2024). Another scenario is that the signal we observe is due to an additional planet interior to the orbit of PDS 70 b. To distinguish between these many scenarios, follow-up observations, at similar wavelengths, will be necessary with the James Webb Interferometer. Analysing follow-up data using an approach similar to Thompson et al. (2023a), using Octofitter555https://sefffal.github.io/Octofitter.jl/dev/ (Thompson et al., 2023b), would allow for a robust detection of any orbital motion, even with low significance detections or non-detections at individual epochs.
Additionally, it is clear from our SNR maps (Figure 5), and our contrast upper limit of 7.56 mags derived in Section 4, that we do not detect a signal (to the north-west of PDS 70 A, at a P.A. 291.8∘) consistent with the point-like feature (PLF) seen by Mesa et al. (2019)/PDS 70 d seen by Christiaens et al. (2024), detected at shorter wavelengths. This indicates that if the signal is a planet, it likely traces bright scattered light emission from a potential planetary envelope instead of indicating another warm source like both PDS 70 b and c. This scenario would be similar to the protoplanet candidate HD 169142 b (Hammond et al., 2023).
Our contrast upper limits, shown in Figure 6, are the deepest limits on additional point source emission within 250 mas in the disk gap of PDS 70, at 4.8 m. This clearly demonstrates the power of the James Webb Interferometer at probing small angular scales compared with direct imaging that achieves significantly better contrasts at larger separations, as is seen by Christiaens et al. (2024).
@@ 6 Conclusions
Furthermore, our results show that NIRISS/AMI can reliably measure relative astrometry and contrasts of young planets in a part of parameter space (small separations and moderate to high contrasts) that is unique to this observing mode, and inaccessible to all other present facilities at 4.8 m. We demonstrate a NIRISS/AMI observing strategy for targets faint enough to acquire greater than 10 groups up the ramp before the signal in the central 9 pixels reaches a total (linearized) value of 30,000 DN (48,000 electrons). We show that by using this stringent data selection criteria, we achieve nearly photon noise limited performance. For observing brighter targets, and to overcome the limitation of the sparsity of the NIRISS AMI uv-coverage, methods that are able to analyze the data directly in the image plane (or using the full extent of the Fourier ”splodges”) will be necessary. A promising approach will be forward modelling of the full optical system and detector systematics (e.g., the brighter-fatter effect, 1/f noise) as is made possible with Lux (Desdoigts et al., 2023).
@@
D.B. and D.J. acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). M.D.F. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303911. J.S.-B. acknowledges the support received from the UNAM PAPIIT project IA 105023.
@@ A.1 Disk Parameters
Table 2 shows the derived disk parameters for the model that used uniform priors on the planet locations, as well as the model that used priors from Wang et al. (2021b). In both cases, we used Gaussian priors on the disk geometry, with the mean values taken from Keppler et al. (2018) and Keppler et al. (2019) and sensibly small standard deviation values chosen so as to break degeneracies between the disk model parameters, using the previously measured geometry of the disk. This was done because we are not primarily interested in independently measuring the parameters of the disk but wish to produce a sensible model of the disk emission so as to measure the emission from the known planets.
The derived disk parameters are nearly identical between the models using the two sets of planet location priors. This confirms what we see in the correlation plot in Figure 4, that there are no significant correlations between the disk parameters and the planet parameters that are biasing our results. If strong correlations were present we would see significant differences between the model disk parameters between the two cases.
A.2 Planet Parameters
In Table 3 we show the planet parameters, calculated using nested sampling, for both the model that used uniform priors on the locations of PDS 70 b and c, and the model that used the GRAVITY priors on the planet locations. The estimated systematic uncertainties in the contrast, separation and position angle, calculated in Section 3, are not included in the errors shown in Table 3.
A.3 Image Plane Residuals
In Figure 9 we plot the deconvolved image plane residuals from subtracting off different components of our PDS 70, disk, and planets model fit to the squared visibilities and closure phases. The deconvolved image plane residuals are primarily shown to verify the fitting to the interferometric observables and are not directly comparable to PSF subtracted direct images, due to the possibility of complex residual features introduced by the highly structured AMI PSF and the deconvolution process.
As also noted in the main text, we tentatively observe compact emission at a similar position to the mm emission seen by Balsalobre-Ruza et al. (2023) (top left panel of Figure 9.). The tentative compact emission seen in the image plane residuals is at the same position angle and is consistent with the signal that we see in the residual closure phases, shown in Figure 5. We note that the deconvolved image plane residuals are a poor diagnostic for measuring the noise in AMI data, because much of the residual signal is caused by differential piston errors (the dominant phase error term on the time scale of hours) to which the closure phase is immune. Thus, the SNR plots shown in Figure 5 give a much clearer picture of the level of signal left in the data after model removal of the disk and two planets, which gives us confidence in the detection of the compact emission to the south of PDS 70 A, as this cannot be caused by differential piston errors.
A.4 SED Fit Parameters
| Parameters | Prior | Uniform prior | GRAVITY prior |
| FWHMr (as) | |||
| P.A. (∘) | |||
@@ A.4 SED Fit Parameters
| Sep. Prior | P.A. Prior | Separation | Position Angle | Contrast | Flux | SNR | 2 | |
| (mas) | (∘) | (mas) | (∘) | (mag) | (W m-2 m-1) | |||
| PDS 70 b | 21.1 | -30.0 | ||||||
| 27.9 | -32.9 | |||||||
| PDS 70 c | 10.6 | |||||||
| 11.9 |
Notes: The median along with the 16th and 84th percentiles of the marginalized posteriors are reported. denotes a uniform prior with lower and upper bounds a and b, respectively. denotes a Gaussian prior, with mean and standard deviation .
1 From the orbit posterior published in (Wang et al., 2021b), calculated using whereistheplanet (Wang et al., 2021a).
2 is the maximum likelihood of the full model including the star, both planets and the disk.
@@ Appendix B The Brighter-Fatter Effect
In Figure 10, we show the evolution of the pixel intensities for the central 9 pixels of the PSF, for PDS 70 and HD 123991. We show how the linearized ramp data evolves (blue), as each pixel accumulates signal, along with how this affects the calibrated rate data (red), output by the jwst pipeline (the calints data). We note that for the central pixel, the detected signal decreases, apparently at all signal levels, with the signal in most of the adjacent pixels either rising or falling, with noticeable differences between PDS 70 and HD 123991. The chosen cutoff value, where we discard all data above this group, is denoted by the vertical dashed red line. For all pixels other than the central pixel, there is no significant signal deficit/excess up to this cutoff value.

### v2

@@ 5.2 Contrast Limits and the Nature of the Signal Seen in the Residuals
The residual signal shown in Figure 7, is in a direction significantly offset from the expected forward scattering direction of the inner disk. This indicates that what we observe is not due to a simple inner disk structure, and may hint at a complex inner disk morphology such as a spiral or clumpy features, as has been suggested by Casassus & Cárcamo (2022), from ALMA observations in the mm, and by Mesa et al. (2019) from near-infrared observations with VLT/SPHERE. It is also possible that the asymmetry we observe is related to the larger separation, 180 mas, ALMA compact emission seen by Balsalobre-Ruza et al. (2023), as it is at very similar position angle. It might even highlight extended emission from an accretion stream between PDS 70 b and c, tentatively detected by Christiaens et al. (2024). Another scenario is that the signal we observe is due to an additional planet interior to the orbit of PDS 70 b. To distinguish between these many scenarios, follow-up observations, at similar wavelengths, will be necessary with NIRISS/AMI. Analysing follow-up data using an approach similar to Thompson et al. (2023a), using Octofitter555https://sefffal.github.io/Octofitter.jl/dev/ (Thompson et al., 2023b), would allow for a robust detection of any orbital motion, even with low significance detections or non-detections at individual epochs.
Additionally, it is clear from our SNR maps (Figure 7), and our contrast upper limit of 7.04 mags derived in Section 4, that we do not detect a signal (to the north-west of PDS 70 A, at a P.A. 291.8∘) consistent with the point-like feature (PLF) seen by Mesa et al. (2019)/PDS 70 d seen by Christiaens et al. (2024), detected at shorter wavelengths. This indicates that if the signal is a planet, it likely traces bright scattered light emission from a potential planetary envelope instead of indicating another warm source like both PDS 70 b and c. This scenario would be similar to the protoplanet candidate HD 169142 b (Hammond et al., 2023).
Our contrast upper limits, shown in Figure 8, are the deepest limits on additional point source emission within 250 mas in the disk gap of PDS 70, at 4.8 m. This clearly demonstrates the power of NIRISS/AMI at probing small angular scales compared with direct imaging that achieves significantly better contrasts at larger separations, as is seen by Christiaens et al. (2024) (see also Carter et al., 2023).
@@ 6 Conclusions
Furthermore, our results show that NIRISS/AMI can reliably measure relative astrometry and contrasts of young planets in a part of parameter space (small separations and moderate to high contrasts) that is unique to this observing mode, and inaccessible to all other present facilities at 4.8 m. We demonstrate a NIRISS/AMI observing strategy for targets faint enough to acquire greater than 10 groups up the ramp before the signal in the central 9 pixels reaches a total (linearized) value of 30,000 DN (48,000 electrons). We show that by using this stringent data selection criteria, we achieve nearly photon noise limited performance. For observing brighter targets, and to overcome the limitation of the sparsity of the NIRISS AMI uv-coverage, methods that are able to analyze the data directly in the image plane (or using the full extent of the Fourier “splodges”) will be necessary. A promising approach will be forward modelling of the full optical system and detector systematics (e.g., the brighter-fatter effect, 1/f noise) as is made possible with Lux (Desdoigts et al., 2023).
@@
The authors would like to thank the referee for their comments and suggestions that significantly strengthened the paper. The authors also thank Gregory Herczeg for his helpful comments.
@@
D.B. and D.J. acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). G.C. thanks the Swiss National Science Foundation for financial support under grant numbers P500PT_206785 and P5R5PT_225479. M.D.F. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303911. J.S.-B. acknowledges the support received from the UNAM PAPIIT project IA 105023.
@@ A.1 Disk Parameters
Table 2 shows the derived disk parameters for each of the Henyey-Greenstein, Gaussian and power law models. In all cases, we used Gaussian priors on the disk geometry, with the mean values taken from Keppler et al. (2018) and Keppler et al. (2019) and sensibly small standard deviation values chosen so as to break degeneracies between the disk model parameters, using the previously measured geometry of the disk. This was done because we are not primarily interested in independently measuring the parameters of the disk but wish to produce a sensible model of the disk emission so as to measure the emission from the known planets.
The derived disk parameters are nearly identical between the models. This confirms what we see in the correlation plot in Figure 5, that there are no significant correlations between the disk parameters and the planet parameters that could be biasing our results. If strong correlations were present we would see significant differences between the model disk parameters between the two cases.
| Parameters | Prior | Henyey-Greenstein | Gaussian | Power Law |
| FWHMr (as) | ||||
| P.A. (∘) | ||||
| log | ||||
@@
4 The logarithm of the Bayesian evidence for each model.
A.2 Planet Parameters
In Table 3 we show the planet parameters, calculated using nested sampling, from each of the Henyey-Greenstein, Gaussian and power law models.
A.3 SED Fit Parameters
| PDS 70 b | PDS 70 b | |
| Henyey-Greenstein | ||
| Separation (mas) | ||
| Position Angle (∘) | ||
| Contrast (mag) | ||
| Gaussian | ||
| Separation (mas) | ||
| Position Angle (∘) | ||
| Contrast (mag) | ||
| Power Law | ||
| Separation (mas) | ||
| Position Angle (∘) | ||
| Contrast (mag) |
Notes: We report the median along with the 16th and 84th percentiles of the marginalized posteriors of each model.
@@ Appendix B The Brighter-Fatter Effect
In Figure 11, we show the evolution of the pixel intensities for the central 9 pixels of the PSF, for PDS 70 and HD 123991. We show how the linearized ramp data evolves (orange), as each pixel accumulates signal, along with how this affects the calibrated rate data (red), output by the jwst pipeline (the calints data). We note that for the central pixel, the detected signal decreases, apparently at all signal levels, with the signal in most of the adjacent pixels either rising or falling, with noticeable differences between PDS 70 and HD 123991. The chosen cutoff value, where we discard all data above this group, is denoted by the vertical dashed red line. For all pixels other than the central pixel, there is no significant signal deficit/excess up to this cutoff value.